# Backdoor Attack in Prompt-Based Continual Learning

### v1

@@ 3.5 Towards an Authentic Backdoor Trigger
Are we truly optimizing a backdoor trigger? As discussed in Section 3.2, optimizing the trigger with these objectives can unintentionally transform it into adversarial noise. While our static-dynamic framework can generate a robust trigger that withstands intense incremental learning processes, it might deviate into adversarial perturbation. To further explore this phenomenon, we analyze the output scores in Figure 1(c). The visualization reveals that even when processed by a clean model unaffected by backdoor attacks, the poisoned samples are consistently misclassified towards the target class with dominant scores. This observation prompts a reconsideration of the backdoor trigger optimization process. We discovered that the overconfident score bias towards the target class is primarily induced by the commonly used softmax with cross-entropy loss function. Softmax introduces competition between classes, and the subsequent cross-entropy loss tends to elevate the scores of the target class significantly above the others. This pronounced bias compels the trigger to act like adversarial noise.
Sigmoid with binary cross entropy loss To reduce biases, we mitigate the competition between the target class and other classes caused by the relative scoring of softmax by employing a sigmoid function after the logits to compute output scores. This approach shifts the optimization focus towards independently increasing the scores of target classes rather than suppressing others. Subsequently, we utilize binary cross-entropy loss to enable independent optimization processes. Following [8], the gradient of the loss at score $(s_{j})$ for class $j$ is computed as $\frac{\partial\mathcal{L}_{\mathrm{BCE}}(\boldsymbol{\theta})}{\partial s_{j}}=\sigma(s_{j})-\mathbb{I}\{j=\tilde{y}\}$, thereby constraining the score of the target class to a certain level regardless of the scores of other classes. As a result, during inference with a non-backdoored clean model, the output scores are more balanced between classes, as shown in Figure 1(f). This balance prevents the problem of generating adversarial noise when optimizing the backdoor trigger.
@@ 4 Experiments
In this section, we first describe the experimental setups, followed by presenting the results in four key aspects: the overall backdooring ability of AOP, its performance with different surrogate datasets, the robustness of AOP with varying attack times, and the efficacy of adopting BCE in preventing the generation of adversarial perturbations. Further discussions on performance, visualizations, baselines, efficacy against defenses, and poisoning rate sensitivity are deferred to Appendix D.
@@ 4.1 Experimental Setup
Victim incremental learners
We evaluate our continual backdoor framework against 6 prompt-based continual learning methods: L2P [58], DualPrompt [57], HiDe-Prompt [55], CODA-Prompt [45], and two variants of PGP [40], namely L2P-PGP and DualPrompt-PGP. We follow the original settings and implementations of each method. All learners utilize the ViT-B/16 backbone [14], pre-trained on ImageNet-1K [41], except for HiDe-Prompt, which is pre-trained on iBOT-1K [66]. Detailed experimental information is in Appendix C.
@@
Datasets For the victim’s training dataset, we use three variants of ImageNet-R [17]: 5-Split, 10-Split, and 20-Split ImageNet-R. These variants divide the 200 classes of the original dataset into 5, 10, and 20 tasks, respectively. Additionally, we conduct experiments on the 5-Split-CUB200 dataset, which partitions the original CUB200 [52] dataset into 5 tasks, each containing 40 classes. For the attacker’s surrogate dataset, we primarily use TinyImageNet [23] for all experiments and CIFAR100 [22] in specific settings.
Backdoor setting Following the guidelines of [64], we set the maximum poison ratio to 25 images, corresponding to $0.1\%$ of ImageNet-R and $0.5\%$ of CUB200. Additionally, we set the upper bound of the $\ell_{\infty}$-norm of triggers to $\frac{16}{255}$, in line with standard practices in the literature [48, 42]. During inference, the trigger is amplified by a factor of $3$ [48, 64].
Metrics The evaluation of our framework utilizes two key metrics: (1) average accuracy (ACC) and (2) attack success rate (ASR). ACC assesses the accuracy of the backdoored model on benign test samples, whereas ASR measures the proportion of attacked samples that the compromised model predicts as the target label, reflecting the backdoor attack’s effectiveness. In the context of continual learning, ACC and ASR at a given time $t$ are averaged across the corresponding metrics for all data from task $1$ to task $t$. All results are averaged over 3 runs for fair comparisons.
@@ 4.2 Effectiveness of AOP
While prompt-based methods share a common core of utilizing prompt pools and selecting relevant prompts for each task or class, each exhibits distinct characteristics. Our framework observes a significantly lower ASR when backdooring CODA-Prompt. This is because CODA-Prompt utilizes all prompts in the prompt pool through its weighted mechanism instead of selecting only the top-K relevant prompts. Consequently, even with triggered samples, clean prompts still exert some influence, leading to degradation in ASR.
@@
Different surrogate datasets Another factor that makes prompt-based continual learning vulnerable is the utilization of prompting. As shown in Figures 1(d) and 1(e), AOP’s triggered samples consistently have the highest similarity with prompt ID 2, which, in contrast, shows the smallest similarity with benign samples. Thus, as discussed in Section 3.3, prompting allows for actual data differences when choosing surrogate datasets. We report the backdoor performance using TinyImageNet and CIFAR100 as surrogate datasets in Table 1. The experiments show consistently high ASR results for both surrogate data choices, confirming the transferability of our continual backdoor framework.
Different attack times We report the ASR in Table 3, considering scenarios where the target class belongs to different tasks that arrive at different times. We observe slight increases in ASR when the attack class is part of later tasks, as it experiences less forgetting. Nonetheless, our method AOP consistently maintains a high ASR, exceeding $99\%$ at all three reported attack times. This convincingly demonstrates that the backdoor knowledge can be effectively transferred to both previously learned and incoming future classes.
Different dynamic rounds We illustrate the attack performance across varying numbers of dynamic rounds in Figure 3. As discussed above, the ASR decreases when tested on the 20-Split-ImageNet. We observe that increasing the number of dynamic rounds does not consistently lead to higher performance. However, from a positive perspective, since the adversary lacks information about the total number of tasks, decreasing and increasing dynamic rounds should not have too much impact on ASR. We emphasize that in long sequence tasks, both ASR and ACC degrade due to forgetting.
Enhancing backdoor authenticity via sigmoid BCE As shown in Table 4, triggers optimized with softmax CE retain considerable scores even when tested on non-backdoored models. This suggests that CE optimization might lead to the generation of adversarial perturbations. Conversely, when optimized using sigmoid BCE, the ASR on clean models remains consistently low. This confirms that adopting BCE can enhance the authenticity of backdoor triggers and avoid generating adversarial noise.
@@ D.1 Further discussion on AOP
In this Appendix, we discuss the differences in ASR when using AOP to backdoor prompt-based continual learners. As shown in Table 2, in most experiments, L2P and PGP achieve the highest ASR, followed by DualPrompt and DualPrompt-PGP.
@@ D.2 Additional comparison between AOP and baseline
In Table 1, we compare AOP and Narcissus, showing that Narcissus experiences catastrophic forgetting. To provide further discussion, in Figure 5, we visualize the ASR flow for each task between our AOP and Narcissus. We trained Narcissus using the same dataset and the same number of epochs as in stages (1) and (2) of our AOP. As visualized in Figure 5, although Narcissus initially achieves high performance, it tends to experience catastrophic forgetting over time. Consequently, the performance gap between AOP and Narcissus increases as the training process continues.
D.3 Defenses
Neural Cleanse
STRIP
FST
We evaluated AOP against a robust fine-tuning-based defense method, FST [35]. FST operates by storing a small amount of clean data to fine-tune the model, reinitializing the classifier weights, and encouraging deviation from the originally compromised weights. We report the performance of FST with respect to different fine-tune data ratios as in the original paper ($2\%$ and $5\%$) and varying weights on the deviation regularizer.
Discussion on potential defenses
As observed in Figures 1(b) and 1(b), poisoned samples consistently exhibit queries for specific prompt IDs, while clean samples demonstrate a more balanced distribution in prompt frequency selection. Consequently, potential defenses against AOP may involve monitoring the frequency selections of test samples during inference. A backdoor flag can be raised if biases in prompt selection frequencies are observed in suspected input samples. Furthermore, drawing inspiration from Fine-Pruning techniques [44], which prune inactive neurons when predicting clean images, one could extend this approach to Prompt-Pruning, effectively eliminating inactive prompts.
Table 5: ACC and ASR of AOP on L2P with 10-Split-ImageNet-R when applying FST as the defense method. Here, $\alpha$ represents the weight of the feature shifting regularization, and $N$ denotes the number of samples saved for finetuning.
@@ D.3 Defenses
$\alpha=2e-5$#\## epochs = 10 $41.78$ $0.00$ $56.87$ $0.00$

### v2

@@ 3.4 Towards an Authentic Backdoor Trigger
Are we truly optimizing a backdoor trigger? While our static-dynamic framework can generate a robust trigger that withstands intense incremental learning processes, it can unintentionally transform into adversarial noise. To further explore this phenomenon, we analyze the output scores in Figure 2(c). The visualization reveals that even when processed by a clean model unaffected by backdoor attacks, the poisoned samples are consistently misclassified towards the target class with dominant scores. This observation prompts a reconsideration of the backdoor trigger optimization process. We discovered that the overconfident score bias towards the target class is primarily induced by the commonly used softmax with cross-entropy loss function. Softmax introduces competition between classes, and the subsequent cross-entropy loss tends to elevate the scores of the target class significantly above the others. This pronounced bias compels the trigger to act like adversarial noise.
Sigmoid with binary cross entropy loss. To reduce biases, we mitigate the competition between the target class and other classes caused by the relative scoring of softmax by employing a sigmoid function after the logits to compute output scores. This approach shifts the optimization focus towards independently increasing the scores of target classes rather than suppressing others. Subsequently, we utilize binary cross-entropy loss to enable independent optimization processes. Following [8], the gradient of the loss at score $(s_{j})$ for class $j$ is computed as $\frac{\partial\mathcal{L}_{\mathrm{BCE}}(\boldsymbol{\theta})}{\partial s_{j}}=\sigma(s_{j})-\mathbb{I}\{j=\tilde{y}\}$, thereby constraining the score of the target class to a certain level regardless of the scores of other classes. As a result, during inference with a non-backdoored clean model, the output scores are more balanced between classes, as shown in Figure 2(f). This balance prevents the problem of generating adversarial noise when optimizing the backdoor trigger.
@@ 4 Experiments
In this section, we first describe the experimental setup, then present the results in five key aspects: the overall backdooring ability of AOP, its performance with different surrogate datasets, the robustness of AOP with varying attack times, the efficacy of adopting BCE in preventing the generation of adversarial perturbations and its effectiveness compared to baselines. Further discussions on performance, visualizations, baselines, efficacy against defenses, and sensitivity to poisoning rates are deferred to the supplementary material.
@@ 4.1 Experimental Setup
Victim incremental learners. We evaluate our continual backdoor framework against 6 prompt-based continual learning methods: L2P [58], DualPrompt [57], HiDe-Prompt [55], CODA-Prompt [45], and two variants of PGP [40], namely L2P-PGP and DualPrompt-PGP. We follow the original settings and implementations of each method. All learners utilize the ViT-B/16 backbone [14], pre-trained on ImageNet-1K [41], except for HiDe-Prompt, which is pre-trained on iBOT-1K [66]. Detailed experimental information is in the supplementary material.
@@
Datasets. For the victim’s training dataset, we follow existing prompt-based continual learning methods [55, 40] and use three variants of ImageNet-R [17]: 5-Split, 10-Split, and 20-Split ImageNet-R. These variants divide the 200 classes of the original dataset into 5, 10, and 20 tasks, respectively. Additionally, we conduct experiments on the 5-Split-CUB200 dataset, which partitions the original CUB200 [52] dataset into 5 tasks, each containing 40 classes. For the attacker’s surrogate dataset, we primarily use TinyImageNet [23] for all experiments and CIFAR100 [22] in specific settings.
Backdoor setting. Following the guidelines of [64], we set the maximum poison ratio to 25 images, corresponding to $0.1\%$ of ImageNet-R and $0.5\%$ of CUB200. Additionally, we set the upper bound of the $\ell_{\infty}$-norm of triggers to $\frac{16}{255}$, in line with standard practices in the literature [48, 42]. During inference, the trigger is amplified by a factor of $3$ [48, 64].
Metrics. The evaluation of our framework utilizes two key metrics: (1) average accuracy (ACC) and (2) attack success rate (ASR). ACC assesses the accuracy of the backdoored model on benign test samples, whereas ASR measures the proportion of attacked samples that the compromised model predicts as the target label, reflecting the backdoor attack’s effectiveness. In the context of continual learning, ACC and ASR at a given time $t$ are averaged across the corresponding metrics for all data from task $1$ to task $t$. All results are averaged over 3 runs for fair comparisons.
@@ 4.2 Effectiveness of AOP
While prompt-based methods share a common core of utilizing prompt pools and selecting relevant prompts for each task or class, each exhibits distinct characteristics. AOP observes a significantly lower ASR when backdooring CODA-Prompt. This is because CODA-Prompt utilizes all prompts in the prompt pool through its weighted mechanism instead of selecting only the top-K relevant prompts. Consequently, even with triggered samples, clean prompts still exert some influence, leading to degradation in ASR.
@@
Different surrogate datasets. Another factor that makes prompt-based continual learning vulnerable is the utilization of prompting. As shown in Figures 2(d) and 2(e), AOP’s triggered samples consistently have the highest similarity with prompt ID 2, which, in contrast, shows the smallest similarity with benign samples. Thus, as discussed in Section 3.2, prompting allows for actual data differences when choosing surrogate datasets. We report the backdoor performance using TinyImageNet and CIFAR100 as surrogate datasets in Table 1. The experiments show consistently high ASR results for both surrogate data choices, confirming the transferability of our continual backdoor framework.
Different attack times. We report the ASR in Table 3, considering scenarios where the target class belongs to different tasks that arrive at different times. We observe slight increases in ASR when the attack class is part of later tasks, as it experiences less forgetting. Nonetheless, our method AOP consistently maintains a high ASR, exceeding $99\%$ at all three reported attack times. This convincingly demonstrates that the backdoor knowledge can be effectively transferred to both previously learned and incoming future classes.
Different dynamic rounds. We illustrate the attack performance across varying numbers of dynamic rounds in Figure 4. As discussed above, the ASR decreases when tested on the 20-Split-ImageNet. We observe that increasing the number of dynamic rounds does not consistently lead to higher performance. However, from a positive perspective, since the adversary is unaware of the total tasks, adjusting the number of dynamic rounds should minimally impact ASR. We emphasize that in long sequence tasks, both ASR and ACC degrade due to forgetting.
Enhancing backdoor authenticity via sigmoid BCE. As shown in Table 4, triggers optimized with softmax CE retain considerable scores even when tested on non-backdoored models. This suggests that CE optimization might lead to the generation of adversarial perturbations. Conversely, when optimized using sigmoid BCE, the ASR on clean models remains consistently low. This confirms that adopting BCE can enhance the authenticity of backdoor triggers and avoid generating adversarial noise.
@@
Comparisons to baselines. We compare the performance of AOP in executing backdoor attacks on prompt-based continual learners against the black-box and clean-label backdoor frameworks, including Narcissus [64] and Label Consistent (LC) [48]. As illustrated in Figure 5, during the first task, both AOP and Narcissus significantly outperform LC, as LC is not designed for data-constrained scenarios. However, as the incremental learning process progresses, Narcissus suffers from forgetting, leading to a sharp decline in its average ASR. In contrast, AOP maintains its backdoor effectiveness throughout all tasks, thanks to its dynamic training strategy.
Figure 5: Average ASR after each task during the attack on L2P and L2P-PGP on CUB200. The figure compares the performance of LC (Label Consistent) [48], Narcissus [64], and AOP (ours).
@@ D.1 Further discussion on AOP
In this Appendix, we discuss the differences in ASR when using AOP to backdoor prompt-based continual learners. As shown in Table 2 of the main manuscript, in most experiments, L2P and PGP achieve the highest ASR, followed by DualPrompt and DualPrompt-PGP.
@@ D.2 Additional comparison between AOP and baseline
In Table 1 of the main manuscript, we compare AOP and Narcissus, showing that Narcissus experiences catastrophic forgetting. To provide further discussion, in Figure 7, we visualize the ASR flow for each task between our AOP and Narcissus. We trained Narcissus using the same dataset and the same number of epochs as in stages (1) and (2) of our AOP. As visualized in Figure 7, although Narcissus initially achieves high performance, it tends to experience catastrophic forgetting over time. Consequently, the performance gap between AOP and Narcissus increases as the training process continues.
Table 5: ACC and ASR of AOP on L2P with 10-Split-ImageNet-R when applying Vanilla FT and FST as the defense methods. Here, $\alpha$ represents the weight of the feature shifting regularization, and $N$ denotes the number of samples saved for finetuning.
@@
Vanilla FT ##\## epochs = 10 $61.25$ $99.55$ $67.24$ $98.96$
##\## epochs = 20 $65.56$ $99.07$ $66.64$ $97.02$
FST $\alpha=2e-5$#\## epochs = 10 $41.78$ $0.00$ $56.87$ $0.00$
@@
D.3 Defenses
Neural Cleanse
STRIP
Feature Shift Tuning
We evaluated AOP against a robust fine-tuning-based defense method, FST [35] in Table 5. FST operates by storing a small amount of clean data to fine-tune the model, reinitializing the classifier weights, and encouraging deviation from the originally compromised weights. We report the performance of FST with respect to different fine-tune data ratios as in the original paper ($2\%$ and $5\%$) and varying weights on the deviation regularizer.
We also evaluate Vanilla Fine-tuning (Vanilla FT) as a potential defense against AOP. As shown in Table 5, while Vanilla FT preserves accuracy on clean samples, it fails to effectively mitigate the impact of AOP.
Discussion on potential defenses
As observed in Figures 2a and 2b, poisoned samples consistently exhibit queries for specific prompt IDs, while clean samples demonstrate a more balanced distribution in prompt frequency selection. Consequently, potential defenses against AOP may involve monitoring the frequency selections of test samples during inference. A backdoor flag can be raised if biases in prompt selection frequencies are observed in suspected input samples. Furthermore, drawing inspiration from Fine-Pruning techniques [44], which prune inactive neurons when predicting clean images, one could extend this approach to Prompt-Pruning, effectively eliminating inactive prompts.